# Fairness and Bias in Algorithmic Hiring: a Multidisciplinary Survey

### v2

@@ 6.2. Visual datasets
Dataset name
Used in
Type
Language
Geography
Target variable
Sensitive variable*
Hiring stage
Chinese Bios
(Zhang et al., 2022)
Textual
Chinese
Synthetic
Mention of job-specific skills in bio
Gender (B, A)
Sourcing
Bias in Bios
(De-Arteaga et al., 2019; Hemamou and Coleman, 2022)
Textual
English
World
Candidate occupation
Gender (B, A)
Sourcing
Engineers & Scientists
(Cowgill, 2018)
Textual
English
Unknown
Candidate received an offer
Unknown
Screening
IT Resumes
(Parasurama et al., 2022; Parasurama and Sedoc, 2021)
Textual
English
US
Candidate was called back
Gender (B, S)
Sourcing
CVs from Singapore
(Deshpande et al., 2020)
Textual
English
Singapore
CV match with job description
Ethnicity (A)
Sourcing
DPG Resumes
(Rus et al., 2022)
Textual
English, Dutch
Netherlands
Candidate industry of interest
Gender (B, A)
Sourcing
ChaLearn First Impression
(Hemamou et al., 2021; Yan et al., 2020; Köchling et al., 2021)
Visual
English
Unknown
Speaker hireability and personality annotated by AMT workers
Gender (B, A), ethnicity (A)
Screening
Student Interviews
(Booth et al., 2021)
Visual
English
Unknown
Speaker hireability annotated by research assistants
Gender (NB, S)
Screening
SHL Interviews
(Singhania et al., 2020)
Visual
English
US, UK, India
Speaker communication skills rated by experts
Gender (B), age, race, country of residence
Screening
Oil Company Interviews
(Kappen and Naber, 2021)
Visual
Dutch
Netherlands
Candidate self-reported commitment
Gender, age
Screening
FairCVs
(Peña et al., 2020; Hemamou and Coleman, 2022)
Visual
English
Synthetic
Synthetic score
Gender (B, A), race (A)
Sourcing
Requirements and Candidates
(Markert et al., 2022)
Tabular
Synthetic
Synthetic
Synthetic score
Synthetic
Sourcing
Jobs and Candidates
(Arafan et al., 2022)
Tabular
Unknown
Netherlands
Candidate recruited or shortlisted
Gender (B, S)
Sourcing
Pymetrics Bias Group
(Wilson et al., 2021)
Tabular
Unknown
Unknown
Similarity to current employees
Gender (S), race (S)
Screening
IBM HR Analytics
(Ghazimatin et al., 2022)
Tabular
Unknown
Unknown
Employee resignation
Gender (B), age, marital status
Evaluation
Resume Search Engines
(Chen et al., 2018a)
Tabular
English
US
Unknown
Gender (B, S)
Sourcing
Walmart Employees
(Burke et al., 2021)
Tabular
English
US
Employee tenure and performance ratings
Synthetic
Screening
Web Developers Field Study
(Avery et al., 2023)
Tabular
English
US
Unknown
Gender (B,S), age (S), ethnicity (S)
Screening
Chinese Job Recommendations
(Zhang and Kuhn, 2022)
Tabular
Chinese
China
Unknown
Gender (B,A), age (A)
Sourcing
Facebook Ads Audiences
(Ali et al., 2019)
Tabular
English
US
User clicks
Gender (B, S), race (B, A)
Sourcing
@@ 7.2. Tasks
Ranking is a common task in the sourcing and screening stages of hiring, performed by search engines or recommender systems. However, the prevalence of classification-based bias metrics in algorithmic hiring (discussed in more detail in Section 4) is reflected in the aforementioned open source packages, which means that measuring and mitigating bias in ranking systems tends to follow the practice of re-purposing classification methods through, e.g., thresholding rankings (i.e., cutting off at $k$). We recommend using fair ranking measures whenever possible for the evaluation of ranking systems, although their adoption can be complicated by the need to model browsing behavior. While several mitigation methods apply to both classification and ranking tasks (e.g. adversarial inference), practitioners should be aware of ranking-specific methods such as DetGreedy.
@@ 7.3. Scalability and efficiency
Technical and infrastructural decisions further affect which bias mitigation methods to consider. For example, Geyik et al. (2019) decouple their post-processing method from specific model choices and properties of input data, which means their method can naturally scale across the different (ranking) systems at LinkedIn. This also means that such a method can be developed and deployed as a standalone component or micro-service, which decreases development time, effort, and alignment, when compared to to pre-processing or in-processing methods that may need to be designed and integrated in complex algorithmic hiring pipelines. In general, post-processing approaches offer an advantage in terms of system integration, but they should be applied with care due to anti-discrimination law (Section 8).
@@ 7.4. Sensitive attribute data availability and usage
The availability of sensitive data is affected by several factors in practice. First, as we identify in Section 8, access to sensitive attributes may be at odds with privacy regulations such as the GDPR, which is an important real-world constraint. Next, availability of and access to sensitive attributes may require job seekers’ explicit consent, which can be difficult to acquire at the scale needed for some bias mitigation methods. Finally, some sensitive attributes, such as age and gender, may have to be recorded in the hiring process for identification purposes, and can hence be assumed to be available for all job seekers. Many other sensitive attributes (e.g., religious beliefs) will not be, and will have to be explicitly requested for bias measurement and mitigation purposes.
This lack of access to sensitive attributes means some bias mitigation methods may be less suited than others; in particular, post-processing methods that require sensitive attributes for all candidates at inference time can be unrealistic. Here, pre-processing bias mitigation methods may prove more useful, as they (only) require access to sensitive attributes at training time, and can operate on attributes even when available for only a subset of job seekers (e.g., those who provided consent). Furthermore, this allows these methods to be deployed and run in isolation from production systems in controlled batch scenarios, which can be another important practical benefit. Examples of these bias mitigation methods used by practitioners include adversarial inference (Rus et al., 2022), and re-balancing training data with synthetic data generation (Arafan et al., 2022).
@@ 7.5. Fairness definitions and intervention targets
In general, the challenge in defining fairness and formulating intervention targets in algorithmic hiring is exacerbated by the multi-stakeholder nature of the hiring domain, where development teams may need to consult legal and compliance teams, HR professionals and recruiters, product managers, and executives, all of whom may need to be informed or provide input. Indeed, these choices should be guided by ethical, social, and legal dimensions. This challenge may explain in part the popularity of disparate impact as a fairness metric, due to its ease of interpretation across a broad and diverse stakeholder group, or the popularity of complying with the EEOC’s 80% rule (Equal Employment Opportunity Commission, 2015) as a fairness target (adopted e.g. by Pymetrics (Wilson et al., 2021)), as it appears, at least superficially, a legally ground target.
@@
In the sourcing or screening stages, depending on the intervention target, outcome fairness may come at the cost of utility; this can happen, for example, when optimizing for gender parity in heavily male or female-dominated industries. However, perhaps surprisingly, different experiments by practitioners show that outcome fairness does not need to come at the cost of utility. First, in their online A/B-testing experiments, Geyik et al. (2019) find their DetGreedy method improves their fairness metric, with no significant impact on utility. In addition, Arafan et al. (2022) find that their pre-processing mitigation method of rebalancing training data may even improve utility over non-bias-mitigated methods, in an offline experiment using real data from an international HR company. Moreover, with respect to impact fairness, Peng et al. (2019) show that “overcompensation” (i.e., artificially over-representing a gender in the output of a ranking system) as an intervention strategy for a hiring algorithm can in some cases mitigate human bias further down the hiring funnel.
Overall, this section describes several practical considerations constraining bias mitigation interventions; desired utility levels are just one dimension, and often not the most restrictive.
@@ 10. Opportunities and Limitations
Bias & validity
consider large candidate pools, reduce human biases, and attract minority candidates
risk of encoding individual biases along with inevitable societal biases; invalid target variables
focus on vulnerable populations beyond acceptance rates; study individual fairness and exploratory policies; carefully scrutinize new technologies
Broader context
trigger positive feedback loops; consider tech-recruiter collaboration
narrow focus on local outcomes can overlook fairness in entire hiring process; risk of repurposing for termination
center on job seeker impacts; identify leaks in pipelines; design for recruiter-tech interaction; beware of performance and tenure prediction
Data
support evaluations of diversity and inclusion
reduced geographical, linguistic, and sensitive attribute coverage
design data collection for diversity; develop IP-friendly audits
Law
apply binding regulation to positively influence industry
legal restrictions on fairness approaches: concerns about discrimination and data protection
multidisciplinary research balancing fairness, privacy, and anti-discrimination; monitor EU AI Act

### v3

@@ 6.2. Visual datasets
Dataset name Used in Type Language Geography Target variable Sensitive variable* Hiring stage
Chinese Bios (Zhang et al., 2022) Textual Chinese Synthetic Mention of job-specific skills in bio Gender (B, A) Sourcing
Bias in Bios (De-Arteaga et al., 2019; Hemamou and Coleman, 2022) Textual English World Candidate occupation Gender (B, A) Sourcing
Engineers & Scientists (Cowgill, 2018) Textual English Unknown Candidate received an offer Unknown Screening
IT Resumes (Parasurama et al., 2022; Parasurama and Sedoc, 2021) Textual English US Candidate was called back Gender (B, S) Sourcing
CVs from Singapore (Deshpande et al., 2020) Textual English Singapore CV match with job description Ethnicity (A) Sourcing
DPG Resumes (Rus et al., 2022) Textual English, Dutch Netherlands Candidate industry of interest Gender (B, A) Sourcing
ChaLearn First Impression (Hemamou et al., 2021; Yan et al., 2020; Köchling et al., 2021) Visual English Unknown Speaker hireability and personality annotated by AMT workers Gender (B, A), ethnicity (A) Screening
Student Interviews (Booth et al., 2021) Visual English Unknown Speaker hireability annotated by research assistants Gender (NB, S) Screening
SHL Interviews (Singhania et al., 2020) Visual English US, UK, India Speaker communication skills rated by experts Gender (B), age, race, country of residence Screening
Oil Company Interviews (Kappen and Naber, 2021) Visual Dutch Netherlands Candidate self-reported commitment Gender, age Screening
FairCVs (Peña et al., 2020; Hemamou and Coleman, 2022) Visual English Synthetic Synthetic score Gender (B, A), race (A) Sourcing
Requirements and Candidates (Markert et al., 2022) Tabular Synthetic Synthetic Synthetic score Synthetic Sourcing
Jobs and Candidates (Arafan et al., 2022) Tabular Unknown Netherlands Candidate recruited or shortlisted Gender (B, S) Sourcing
Pymetrics Bias Group (Wilson et al., 2021) Tabular Unknown Unknown Similarity to current employees Gender (S), race (S) Screening
IBM HR Analytics (Ghazimatin et al., 2022) Tabular Unknown Unknown Employee resignation Gender (B), age, marital status Evaluation
Resume Search Engines (Chen et al., 2018a) Tabular English US Unknown Gender (B, S) Sourcing
Walmart Employees (Burke et al., 2021) Tabular English US Employee tenure and performance ratings Synthetic Screening
Web Developers Field Study (Avery et al., 2023) Tabular English US Unknown Gender (B,S), age (S), ethnicity (S) Screening
Chinese Job Recommendations (Zhang and Kuhn, 2022) Tabular Chinese China Unknown Gender (B,A), age (A) Sourcing
Facebook Ads Audiences (Ali et al., 2019) Tabular English US User clicks Gender (B, S), race (B, A) Sourcing
@@ 7.2. Tasks
Ranking is a common task in the sourcing and screening stages of hiring, performed by search engines or recommender systems. However, the prevalence of classification-based bias metrics in algorithmic hiring (discussed in more detail in Section 4) is reflected in the aforementioned open source packages, which means that measuring and mitigating bias in ranking systems tends to follow the practice of re-purposing classification methods through, e.g., thresholding rankings (i.e., cutting off at $k$). We recommend using fair ranking measures for an evaluation of ranking systems that is more cognizant of user browsing behavior, although, of course, the latter needs to be suitably modeled. While several mitigation methods apply to both classification and ranking tasks (e.g. adversarial inference), practitioners should be aware of ranking-specific methods such as DetGreedy.
@@ 7.3. Scalability and efficiency
Technical and infrastructural decisions further affect which bias mitigation methods to consider. For example, Geyik et al. (2019) decouple their post-processing method from specific model choices and properties of input data, which means their method can naturally scale across the different (ranking) systems at LinkedIn. This also means that such a method can be developed and deployed as a standalone component or micro-service, which decreases development time, effort, and alignment, when compared to pre-processing or in-processing methods that may need to be designed and integrated in complex algorithmic hiring pipelines. In general, post-processing approaches offer an advantage in terms of system integration, but they should be applied with care due to anti-discrimination law (Section 8).
@@ 7.4. Sensitive attribute data availability and usage
The availability of sensitive data is affected by several factors in practice. First, as we identify in Section 8, access to sensitive attributes may be at odds with privacy regulations such as the GDPR, which is an important real-world constraint. Next, availability of and access to sensitive attributes may require job seekers’ explicit consent, which can be difficult to acquire at the scale needed for some bias mitigation methods. Finally, some sensitive attributes, such as age and gender, may have to be recorded in the hiring process for identification purposes, and can hence be assumed to be available for all job seekers. Many other sensitive attributes (e.g., religious beliefs) will not be easily available, and will have to be explicitly requested for bias measurement and mitigation purposes.
This lack of access to sensitive attributes means some bias mitigation methods may be less suited than others, as noted in Section 5.4; in particular, post-processing methods that require sensitive attributes for all candidates at inference time can be unrealistic. Here, pre-processing bias mitigation methods may prove more useful, as they (only) require access to sensitive attributes at training time, and can operate on attributes even when available for only a subset of job seekers (e.g., those who provided consent). Furthermore, this allows these methods to be deployed and run in isolation from production systems in controlled batch scenarios, which can be another important practical benefit. Examples of these bias mitigation methods used by practitioners include adversarial inference (Rus et al., 2022), and re-balancing training data with synthetic data generation (Arafan et al., 2022).
@@ 7.5. Fairness definitions and intervention targets
In general, the challenge in defining fairness and formulating intervention targets in algorithmic hiring is exacerbated by the multi-stakeholder nature of the hiring domain, where development teams may need to consult legal and compliance teams, HR professionals and recruiters, product managers, and executives, all of whom may need to be informed or provide input. Indeed, these choices should be guided by ethical, social, and legal dimensions. This challenge may explain in part the popularity of disparate impact as a fairness metric, due to its ease of interpretation across a broad and diverse stakeholder group, or the popularity of complying with the EEOC’s 80% rule (Equal Employment Opportunity Commission, 2015) as a fairness target (adopted e.g. by Pymetrics (Wilson et al., 2021)), as it appears, at least superficially, a legally grounded target.
@@
In the sourcing or screening stages, depending on the intervention target, outcome fairness may come at the cost of utility; this can happen, for example, when optimizing for gender parity in heavily male or female-dominated industries. However, perhaps surprisingly, different experiments by practitioners show that outcome fairness does not need to come at the cost of utility. First, in their online A/B-testing experiments, Geyik et al. (2019) find their DetGreedy method improves their fairness metric, with no significant impact on utility. In addition, Arafan et al. (2022) find that their pre-processing mitigation method of rebalancing training data may even improve utility over non-bias-mitigated methods, in an offline experiment using real data from an international HR company. Moreover, with respect to impact fairness, Peng et al. (2019) show that “overcompensation” (i.e., artificially over-representing a gender in the output of a ranking system) as an intervention strategy for a hiring algorithm can in some cases mitigate human bias further down the hiring funnel. Overall, this section described several practical considerations constraining bias mitigation interventions; desired utility levels are just one dimension, and often not the most restrictive.
@@ 10. Opportunities and Limitations
Bias & validity consider large candidate pools, reduce human biases, and attract minority candidates risk of encoding individual biases along with inevitable societal biases; invalid target variables focus on vulnerable populations beyond acceptance rates; study individual fairness and exploratory policies; carefully scrutinize new technologies
Broader context trigger positive feedback loops; consider tech-recruiter collaboration narrow focus on local outcomes can overlook fairness in entire hiring process; risk of repurposing for termination center on job seeker impacts; identify leaks in pipelines; design for recruiter-tech interaction; beware of performance and tenure prediction
Data support evaluations of diversity and inclusion reduced geographical, linguistic, and sensitive attribute coverage design data collection for diversity; develop IP-friendly audits
Law apply binding regulation to positively influence industry legal restrictions on fairness approaches: concerns about discrimination and data protection multidisciplinary research balancing fairness, privacy, and anti-discrimination; monitor EU AI Act